# Quantum simulation of many-body dynamics with noise-robust Trotter decomposition based on symmetric structures

### v2

@@ Abstract
The Suzuki-Trotter decomposition, which digitalizes quantum time evolution, provides a promising framework for simulating quantum dynamics on quantum hardware and exploring quantum advantage over classical computation. However, conventional Trotter circuits require a large number of non-local gates, lowering their faithfulness to the ideal dynamics when implemented on current noisy quantum hardware. While most previous studies have focused on circuit optimization, we instead propose a new Trotter decomposition that is intrinsically circuit-efficient for simulating quantum dynamics on near-term devices. Our method substantially reduces the number of CNOT operations compared to conventional Trotter decompositions by exploiting the symmetry of the target model to construct an effective Hamiltonian with fewer two-qubit gates. We demonstrate the noise robustness of the proposed approach through numerical simulations of a nine-site Heisenberg model under realistic noise, and further validate its experimental practicality on the IBM superconducting device, achieving a state fidelity exceeding $0.98$ when combined with quantum error mitigation in the three-site case. The proposed circuit design is also compatible with existing circuit optimization techniques. Our results establish a practical route toward noise-resilient quantum simulation in many-body dynamics.
@@ I Introduction
Through the rapid advance of quantum hardware, quantum simulation has gained emerging attention to simulate physical and chemical models that become practically intractable by classical computational resources Lanes et al. (2025); Huang et al. (2025). Its prominent and versatile applicability lies in non-equilibrium quantum many-body dynamics, where the digital simulation based on Suzuki-Trotter decomposition enables tractable and scalable approximation of real-time evolution Trotter (1959); Suzuki (1976); Hatano and Suzuki (2005). Simulating Trotter steps with quantum computers frees from classical simulation with exponential computational resources, requiring only linear overhead to system size, which is thus seen as one of the applications with potential near-term quantum advantage.
@@
However, the non-negligible noise level and hardware restrictions of current quantum hardware still pose a significant obstacle to the practical realization of such Trotter-based simulations. In particular, superconducting quantum devices ibm (2022), which are among the most extensively developed and commercially accessible platforms, suffer from noisy non-local gates and limited coherence times Mooney et al. (2021); Yang et al. (2022a). Therefore, it is essential to design quantum circuits with reduced depth and fewer non-local gates, such as CNOT gates, to alleviate noise accumulation and improve the fidelity of simulations.
@@
While substantial efforts have focused on optimizing given Trotter circuits under hardware constraints Lötstedt et al. (2023); Chowdhury et al. (2024a); Choi, Chowdhury, and Yu (2025); Yang et al. (2022b); Zhao et al. (2023); Jones et al. (2019), the underlying Trotter decomposition itself sets the fundamental limits of such optimization. In this work, we design a new alternative Trotter decomposition strategy that substantially reduces the number of CNOT gates in use. We exploit the symmetric structure of the given Hamiltonian, particularly, of the $XXX$ Heisenberg model. We transform the three-site Heisenberg Hamiltonian into a more concise two-site effective Hamiltonian through an encoding and decoding procedure, thereby enabling more efficient circuit construction.
@@
This new decomposition with the effective Hamiltonian reduces the average number of CNOT gates in each Trotter step to $2.625$ per qubit, whereas the conventional Trotter circuit requires $3$. The proposed method highlights the potential of reducing the circuit overhead with a more efficient approach for Trotter decomposition rather than merely performing circuit optimization on existing Trotter circuits.
@@
We demonstrate the noise-robustness of our proposed method through numerical simulation. Our method outperforms the conventional Trotter decomposition in simulating the time evolution of the nine-site $XXX$ Heisenberg model under depolarizing noise. We also simulate the time evolution of a three-site $XXX$ Heisenberg model on the superconducting quantum device ibmq_jakarta provided by the IBM Quantum Platform ibm (2022). Using quantum error mitigation (QEM) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Endo, Benjamin, and Li (2018); Giurgica-Tiron et al. (2020); Yang, Raymond, and Uno (2022); Yang et al. (2023); Cai et al. (2023); Endo et al. (2021), we achieve the target state fidelity over $0.98$ on ibmq_jakarta. In implementing the proposed Trotter circuits, our method finds further compatibility with circuit optimization with the Qiskit package Javadi-Abhari et al. (2024). The overall experiments demonstrate that our method offers not only a novel, efficient Trotter decomposition scheme but also a practical and feasible solution for simulating physical models on current quantum hardware.
@@ II Suzuki-Trotter Decomposition
We consider $N$-site $J=1$ XXXXXX Heisenberg Hamiltonian with $N=4M+1$, $M\in\mathbb{Z}_{\geq 0}$, and open boundary condition formalized as
@@
where $\cdot$ denotes the inner product of three components of the Pauli operators of $i$-th site $\{\hat{\sigma}_{x}^{(i)},\hat{\sigma}_{y}^{(i)},\hat{\sigma}_{z}^{(i)}\}$ defined as
@@
Given a Heisenberg Hamiltonian $\hat{H}=\hat{O}_{1}+\hat{O}_{2}$ that consists of two non-commutative operators $\hat{O}_{1}$ and $\hat{O}_{2}$, i.e. $\hat{O}_{1},\hat{O}_{2}]\not=0$, the conventional Trotter decomposition with $n$ steps Trotter (1959); Suzuki (1976); Hatano and Suzuki (2005) approximates the evolution of this Hamiltonian with
$\hat{U}(t)=\left(\exp(-{\mathrm{i}}\hat{O}_{2}\Delta t)\exp(-{\mathrm{i}}\hat{O}_{1}\Delta t)\right)^{n}+\mathcal{O}(Mn^{-1}),$ (3)
@@
$\hat{O}_{1}$ $=\sum_{i=1}^{2M}\vec{\sigma}^{(2i-1)}\cdot\vec{\sigma}^{(2i)},$ (4)
$\hat{O}_{2}$ $=\sum_{i=1}^{2M}\vec{\sigma}^{(2i)}\cdot\vec{\sigma}^{(2i+1)}.$ (5)
@@
$\hat{u}^{(i)}(\Delta t)=\exp(-{\mathrm{i}}\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(i+1)}\Delta t).$ (6)
Since any two-qubit unitary operation can be realized by three CNOT gates Vatan and Williams (2004); Vidal and Dawson (2004); Shende, Bullock, and Markov (2006), Fig. 1 provides a quantum circuit to implement $\hat{u}^{(i)}(t)$.
@@
Using this circuit block, the Trotter decomposition in Eq. (3) is then constructed by the quantum circuit in Fig. 2. Since each time step has two layers of operator $\hat{u}^{(i)}(t)$, the averaged number of CNOT gates applied over each qubit is three, which represents the most efficient circuit construction regarding the CNOT overhead known to date Chowdhury et al. (2024b).
Figure 2: The quantum circuit of the time evolution operator using the Trotter decomposition in Eq. (3) for $N=5$. The operation corresponding to the propagation of each single time step $\Delta t$ is enclosed by a dashed box.
@@ III Proposed Decomposition
In order to reduce the number of CNOT gates in each single time step, we propose a new Trotter decomposition. First, we change the way of partitioning the Heisenberg Hamiltonian into $\hat{H}=\hat{O^{\prime}}_{0}+\hat{O^{\prime}}_{1}+\hat{O^{\prime}}_{2}$ with the operators
$\hat{O}^{\prime}_{0}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)}+\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)},$ (7)
$\hat{O}^{\prime}_{1}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i-3)}\cdot\vec{\sigma}^{(4i-2)},$ (8)
$\hat{O}^{\prime}_{2}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(4i)}\cdot\vec{\sigma}^{(4i+1)}.$ (9)
@@
Using $\hat{O}^{\prime}_{0}$, $\hat{O}^{\prime}_{1}$, and $\hat{O}^{\prime}_{2}$ with $m$ proposed Trotter blocks for evolution time $\Delta t^{\prime}=t/m$ is then defined as
@@
$\begin{split}&\hat{U}_{\mathrm{Heis}}(t)\\ &=\left(\exp(-{\mathrm{i}}\hat{O}^{\prime}_{1}\Delta t^{\prime})\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})\exp(-{\mathrm{i}}\hat{O}^{\prime}_{2}\Delta t^{\prime})\right)^{m}\\ &\quad+\mathcal{O}\left(Mm^{-1}\right).\end{split}$ (10)
@@
Remarkably, the number of required time steps $m$ in Eq. (10) is smaller than $n$ in Eq. (3) for the same residual error of Trotter decomposition. This arises from the fact that Eq. (10) implements the time evolution operator $e^{-{\mathrm{i}}\hat{O^{\prime}}_{0}\Delta t^{\prime}}$ directly instead of decomposing it into $\prod_{i=1}^{M}e^{-{\mathrm{i}}\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)}\Delta t^{\prime}}e^{-{\mathrm{i}}\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)}\Delta t^{\prime}}$ that causes additional error in the order of $\mathcal{O}(Mm^{-1})$. This implies that the residual error of Eq. (10) is reduced to three-quarters of that of the conventional Trotter decomposition for $m=n$. As a result, using $m=\frac{3}{4}n$ steps of the proposed Trotter blocks achieves the same level of residual error as the conventional Trotter blocks.
In designing a quantum circuit for this new Trotter decomposition, we can use $\{\hat{u}^{(4i-1)}(\Delta t^{\prime})\}_{i=1}^{M}$ and $\{\hat{u}^{(4i)}(\Delta t^{\prime})\}_{i=1}^{M}$ to construct the time evolution operators $\exp(-{\mathrm{i}}\hat{O}_{1}^{\prime}\Delta t^{\prime})$ and $\exp(-{\mathrm{i}}\hat{O}_{2}^{\prime}\Delta t^{\prime})$ in Eq. (10) since they share the same form as Eq. (6). Thus, it suffices to design an efficient circuit of the time evolution operator $\exp(-{\mathrm{i}}\hat{O}^{\prime}_{0}\Delta t^{\prime})$, where $\hat{O}_{0}^{\prime}$ takes the form of a three-site $XXX$ Heisenberg Hamiltonian $\hat{H}_{3}=\vec{\sigma}^{(1)}\cdot\vec{\sigma}^{(2)}+\vec{\sigma}^{(2)}\cdot\vec{\sigma}^{(3)}$.
@@
To construct an efficient quantum circuit of $\exp(-{\mathrm{i}}\hat{H}_{3}t)$, we aim to compress it to a smaller subsystem by deriving an effective Hamiltonian $H_{\mathrm{eff}}$. We focus on the fact that $\hat{H}_{3}$ commutes with $\hat{\sigma}_{\mu}^{(1)}\otimes\hat{\sigma}_{\mu}^{(2)}\otimes\hat{\sigma}_{\mu}^{(3)}$, i.e. $\hat{H}_{3},\hat{\sigma}_{\mu}^{(1)}\otimes\hat{\sigma}_{\mu}^{(2)}\otimes\hat{\sigma}_{\mu}^{(3)}]=0$. This yields simultaneous eigenstates as follows:
@@
[…]
@@
$\begin{split}\hat{H}_{\mathrm{eff}}&=\hat{U}_{\mathrm{enc}}\hat{H}_{3}\hat{U}_{\mathrm{enc}}^{\dagger}\\ &=\hat{\sigma}_{x}^{(1)}+\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{z}^{(1)}+\hat{\sigma}_{z}^{(2)}-(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{x}^{(1)}\otimes\hat{\sigma}_{z}^{(2)}).\end{split}$ (15)
This yields the following equivalence between the time evolution operators:
@@
Figure 4: The quantum circuit of the time evolution operator $e^{-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}}$ given by Eq. (17).
@@
To design an efficient circuit implementation of $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$, we further expand $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$ in the following form
@@
$\begin{split}&\exp(-i\hat{H}_{\mathrm{eff}}\Delta t^{\prime})\\ &=e^{-{\mathrm{i}}(\hat{\sigma}_{z}^{(1)}+\hat{\sigma}_{x}^{(2)})\Delta t^{\prime}}e^{{\mathrm{i}}(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{z}^{(2)}\otimes\hat{\sigma}_{x}^{(1)})\Delta t^{\prime}}e^{-{\mathrm{i}}(\hat{\sigma}_{z}^{(2)}+\hat{\sigma}_{x}^{(1)})\Delta t^{\prime}}\\ &\quad+\mathcal{O}(m^{-3}),\end{split}$ (17)
where the residual error decreases faster than that of the conventional Trotter decomposition. This provides a quantum circuit of the time evolution $\exp(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime})$, represented in Fig. 4.
Figure 5: (a) A concise form of the quantum circuit of the time evolution operator, denoting $\Delta t^{\prime}$ time evolution. (b) The quantum circuit of the new notation of the two-qubit unitary gate $\hat{u}^{\prime}$ seen in (a).

### v3

@@ Abstract
The Suzuki-Trotter decomposition, which digitalizes quantum time evolution, provides a promising framework for simulating quantum dynamics on quantum hardware and exploring quantum advantage over classical computation. However, conventional Trotter circuits require a large number of non-local gates, lowering their faithfulness to the ideal dynamics when implemented on current noisy quantum hardware. While most previous studies have focused on circuit optimization, we instead propose a new Trotter decomposition that is intrinsically circuit-efficient for simulating quantum dynamics on near-term devices. Our method substantially reduces both the residual error by Trotter decomposition and the number of CNOT operations compared to conventional Trotter decompositions by exploiting the symmetry of the target model to construct an effective Hamiltonian with fewer two-qubit gates. We demonstrate the noise robustness of the proposed approach through numerical simulations of a nine-site Heisenberg model under realistic noise, and further validate its experimental practicality on the IBM superconducting device, achieving a state fidelity exceeding $0.98$ when combined with quantum error mitigation in the three-site case. The proposed circuit design is also compatible with existing circuit optimization techniques. Our results establish a practical route toward noise-resilient quantum simulation in many-body dynamics.
@@ I Introduction
Through the rapid advance of quantum hardware, quantum simulation has gained emerging attention to simulate physical and chemical models that become practically intractable by classical computational resources Lanes2025A; Huang2025The. Its prominent and versatile applicability lies in non-equilibrium quantum many-body dynamics, where the digital simulation based on Suzuki-Trotter decomposition enables tractable and scalable approximation of real-time evolution Trotter1959On; Suzuki1976Generalized; Hatano2005Finding. Simulating Trotter iterations with quantum computers frees from classical simulation with exponential computational resources, requiring only linear overhead to system size, which is thus seen as one of the applications with potential near-term quantum advantage.
@@
However, the non-negligible noise level and hardware restrictions of current quantum hardware still pose a significant obstacle to the practical realization of such Trotter-based simulations. In particular, superconducting quantum devices ibm_quantum, which are among the most extensively developed and commercially accessible platforms, suffer from noisy non-local gates and limited coherence times mooney2021generation; Yang2022Testing. Therefore, it is essential to design quantum circuits with reduced depth and fewer non-local gates, such as CNOT gates, to alleviate noise accumulation and improve the fidelity of simulations.
@@
While substantial efforts have focused on optimizing given Trotter circuits under hardware constraints Lotstedt2023Error-mitigated; Chowdhury2024Enhancing; Choi2025Quantum; YangPRA2022; ZhaoPRXQ2023; Jones2019ACM, the underlying Trotter decomposition itself sets the fundamental limits of such optimization. In this work, we design a new alternative Trotter decomposition strategy that substantially reduces both the residual error by Trotter decomposition and the number of CNOT gates in use. To achieve this, we exploit the symmetric structure of the given Hamiltonian, particularly, of the $XXX$ Heisenberg model. In particular, we transform the three-site Heisenberg Hamiltonian into a more concise two-site effective Hamiltonian through an encoding and decoding procedure. This transformation enables a faster convergence rate of the Trotter iterations compared with using the conventional Trotter decomposition, thereby requiring fewer Trotter iterations to achieve the same level of residual error.
@@
This new decomposition with the effective Hamiltonian reduces the average number of CNOT gates in each Trotter step to $1.75$ per qubit, whereas the conventional Trotter circuit requires $3$. The proposed method highlights the potential of reducing the circuit overhead with a more efficient approach for Trotter decomposition rather than merely performing circuit optimization on existing Trotter circuits. The schematic illustration of the proposed method can be found in Fig. 1.
@@
We demonstrate the noise-robustness of our proposed method through numerical simulation. Our method outperforms the conventional Trotter decomposition in simulating the time evolution of the nine-site $XXX$ Heisenberg model in both setups without noise and with depolarizing noise. We also simulate the time evolution of a three-site $XXX$ Heisenberg model on the superconducting quantum device ibmq_jakarta provided by the IBM Quantum Platform ibm_quantum. Using quantum error mitigation (QEM) li2017efficient; Temme_2017; endo2018practical; digital_zne; yang2022efficient; Yang2025Resource-Efficient; cai2023quantum; endo2021hybrid, we achieve the target state fidelity over $0.98$ on ibmq_jakarta. In implementing the proposed Trotter circuits, our method finds further compatibility with circuit optimization with the Qiskit package qiskit2024. The overall experiments demonstrate that our method offers not only a novel, efficient Trotter decomposition scheme but also a practical and feasible solution for simulating physical models on current quantum hardware.
Figure 1: The schematic illustration of the proposed framework. In simulating the time evolution of a given Hamiltonian, we exploit its symmetric structure to construct Trotter blocks with a faster convergence rate and fewer CNOT gates based on the effective Hamiltonian.
@@ II Suzuki-Trotter Decomposition
We consider $N$-site $J=1$ XXXXXX Heisenberg Hamiltonian with $N=2M+1$, $M\in\mathbb{Z}_{\geq 0}$, and open boundary condition formalized as
@@
where $\cdot$ denotes the inner product of three components of the Pauli operators of $i$-th site, $\hat{\sigma}_{x}^{(i)}:=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$$\hat{\sigma}_{y}^{(i)}:=\begin{pmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{pmatrix}$, and $\hat{\sigma}_{z}^{(i)}:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$, defined as
@@
Given a Heisenberg Hamiltonian $\hat{H}=\hat{O}_{1}+\hat{O}_{2}$ that consists of two non-commutative operators $\hat{O}_{1}$ and $\hat{O}_{2}$, i.e. $\hat{O}_{1},\hat{O}_{2}]\not=0$, the conventional Trotter decomposition with $n$ steps Trotter1959On; Suzuki1976Generalized; Hatano2005Finding approximates the evolution of this Hamiltonian with
$\begin{split}\hat{U}(t)&=\biggl(\exp\left(-{\mathrm{i}}\hat{O}_{2}\Delta t\right)\exp\left(-{\mathrm{i}}\hat{O}_{1}\Delta t\right)\biggr)^{n}\\ &\quad+\mathcal{O}\left(\hat{\epsilon}_{1}n^{-1}\right),\end{split}$ (3)
@@
$\hat{O}_{1}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(2i-1)}\cdot\vec{\sigma}^{(2i)},$ (4)
$\hat{O}_{2}$ $=\sum_{i=1}^{M}\vec{\sigma}^{(2i)}\cdot\vec{\sigma}^{(2i+1)},$ (5)
and $\hat{\epsilon}_{1}$ is the operator in the error term as follows:
$\hat{\epsilon}_{1}$ $=[\hat{O}_{1},\hat{O}_{2}]$
$=\sum_{i=2}^{N-1}(-1)^{i}[\vec{\sigma}^{(i-1)}\cdot\vec{\sigma}^{(i)},\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(i+1)}].$ (6)
@@
$\hat{u}^{(i)}(\Delta t)=\exp\left(-{\mathrm{i}}\vec{\sigma}^{(i)}\cdot\vec{\sigma}^{(i+1)}\Delta t\right).$ (7)
Since any two-qubit unitary operation can be realized by three CNOT gates Vatan2004Optimal; Vidal2004Universal; Shende2006Synthesis, Fig. 2 provides a quantum circuit to implement $\hat{u}^{(i)}(t)$.
@@
Using this circuit block, the Trotter decomposition in Eq. (3) is then constructed by the quantum circuit in Fig. 3. Since each time step has two layers of operator $\hat{u}^{(i)}(t)$, the averaged number of CNOT gates applied over each qubit is three, which represents the most efficient circuit construction regarding the CNOT overhead known to date PRR2024Talal.
Figure 3: The quantum circuit of the time evolution operator using the Trotter decomposition in Eq. (3) for the 55-site system. The operation corresponding to the propagation of each single time step $\Delta t$ is enclosed by a dashed box.
@@ III Proposed Decomposition
Figure 4: (a) The schematic illustration of the conventional Trotter decomposition that takes non-commutative Trotter blocks with propagator $\hat{u}(\Delta t)$ according to the native edge structure of the given Hamiltonian. (b) The schematic illustration of the proposed Trotter decomposition that takes non-commutative Trotter blocks with propagator $\hat{U}_{\mathrm{enc}}^{\dagger}\exp\left(-\mathrm{i}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)\hat{U}_{\mathrm{enc}}$ according to the hyper-edge of the set with three neighboring vertices in the given Hamiltonian.
@@
In order to reduce the number of CNOT gates in each single time step, we propose a new Trotter decomposition. First, we partition the Heisenberg Hamiltonian into $\hat{H}=\hat{A}_{1}+\hat{B}_{1}$ with the operators
$\hat{A}_{1}$ $=\sum_{i=1}^{M^{\prime}}\vec{\sigma}^{(4i-3)}\cdot\vec{\sigma}^{(4i-2)}+\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)},$ (8)
$\hat{B}_{1}$ $=\sum_{i=1}^{M^{\prime\prime}}\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)}+\vec{\sigma}^{(4i)}\cdot\vec{\sigma}^{(4i+1)},$ (9)
where $M^{\prime}=M^{\prime\prime}=M/2$ for $M\in 2\mathbb{Z}$ and $M^{\prime}=M^{\prime\prime}+1=(M+1)/2$ for $M\not\in 2\mathbb{Z}$. Likewise, we also propose another decomposition $\hat{H}=\hat{A}_{2}+\hat{B}_{2}$ with the operators
$\hat{A}_{2}$ $=\delta_{M,2N^{\prime}+1}\vec{\sigma}^{(2M)}\cdot\vec{\sigma}^{(2M+1)}$
$+\sum_{i=1}^{N^{\prime}}\vec{\sigma}^{(4i-2)}\cdot\vec{\sigma}^{(4i-1)}+\vec{\sigma}^{(4i-1)}\cdot\vec{\sigma}^{(4i)},$ (10)
$\hat{B}_{2}$ $=\vec{\sigma}^{(1)}\cdot\vec{\sigma}^{(2)}+\delta_{M,2N^{\prime}}\vec{\sigma}^{(2M)}\cdot\vec{\sigma}^{(2M+1)}$
$+\sum_{i=1}^{N^{\prime\prime}-1}\vec{\sigma}^{(4i)}\cdot\vec{\sigma}^{(4i+1)}+\vec{\sigma}^{(4i+1)}\cdot\vec{\sigma}^{(4i+2)},$ (11)
where $N^{\prime}=N^{\prime\prime}=M/2$ for $M\in 2\mathbb{Z}$ and $N^{\prime}=N^{\prime\prime}-1=(M-1)/2$ for $M\not\in 2\mathbb{Z}$.
Using $\hat{A}_{k}$ and $\hat{B}_{k}$, we define a single Trotter iteration as a sequence of propagators
$\exp\left(-{\mathrm{i}}\hat{B}_{2}\Delta t^{\prime}\right)\exp\left(-{\mathrm{i}}\hat{A}_{2}\Delta t^{\prime}\right)\exp\left(-{\mathrm{i}}\hat{B}_{1}\Delta t^{\prime}\right)\exp\left(-{\mathrm{i}}\hat{A}_{1}\Delta t^{\prime}\right).$ (12)
@@
Then, the unitary operation that evolves the system for the time $t$ using $m$ proposed Trotter iterations with $\Delta t^{\prime}=t/(2m)$ is described as
@@
$\begin{split}\hat{U}(t)&=\Biggl(\exp\left(-{\mathrm{i}}\hat{B}_{2}\Delta t^{\prime}\right)\exp\left(-{\mathrm{i}}\hat{A}_{2}\Delta t^{\prime}\right)\\ &\quad\quad\quad\exp\left(-{\mathrm{i}}\hat{B}_{1}\Delta t^{\prime}\right)\exp\left(-{\mathrm{i}}\hat{A}_{1}\Delta t^{\prime}\right)\Biggr)^{m}\\ &\quad+\mathcal{O}\left(\hat{\epsilon}_{2}m^{-1}\right),\end{split}$ (13)
with an operator $\hat{\epsilon}_{2}$ defined as
$\hat{\epsilon}_{2}$ $=\frac{1}{4}[\hat{A}_{1},\hat{B}_{1}]+\frac{1}{4}[\hat{A}_{2},\hat{B}_{2}]$
$=\frac{1}{4}\sum_{j=1}^{M}(-1)^{j}[\vec{\sigma}^{(2j-1)}\cdot\vec{\sigma}^{(2j)},\vec{\sigma}^{(2j)}\cdot\vec{\sigma}^{(2j+1)}]$
$\quad+\frac{1}{4}\sum_{j=1}^{M-1}(-1)^{j}[\vec{\sigma}^{(2j)}\cdot\vec{\sigma}^{(2j+1)},\vec{\sigma}^{(2j+1)}\cdot\vec{\sigma}^{(2j+2)}],$ (14)
@@
Remarkably, the number of required Trotter iterations $m$ in Eq. (13) is smaller than $n$ in Eq. (3) to achieve the same level of residual error by Trotter decomposition. This arises from the fact that, as shown in Fig. 4, the proposed Trotter iteration has sparser decomposition interval than the conventional one, which yields fewer combinations of non-commutative operators that contributes to the residual error. In particular, the residual error of Eqs. (13) and  (14) is reduced to one-quarters of that of the conventional Trotter decomposition for $m=n$. This implies that using $m=n/4$ iterations of the proposed Trotter blocks achieves the same level of residual error as the conventional Trotter blocks.
@@
Next, we aim at designing efficient quantum circuits of the time evolution operator $\exp\left(-{\mathrm{i}}\hat{A}_{k}\Delta t^{\prime}\right)$ and $\exp\left(-{\mathrm{i}}\hat{B}_{k}\Delta t^{\prime}\right)$, where $\hat{A}_{k}$ and $\hat{B}_{k}$ take the form of a three-site $XXX$ Heisenberg Hamiltonian $\hat{H}_{3}=\vec{\sigma}^{(1)}\cdot\vec{\sigma}^{(2)}+\vec{\sigma}^{(2)}\cdot\vec{\sigma}^{(3)}$. To construct an efficient quantum circuit of $\exp\left(-{\mathrm{i}}\hat{H}_{3}t\right)$, we compress it to a smaller subsystem by deriving an effective Hamiltonian $H_{\mathrm{eff}}$ focusing on the $\mathrm{SU}(2)$ symmetry that $\hat{H}_{3}$ is equipped with. We use the fact that the three-site Hamiltonian $\hat{H}_{3}$ commutes with the operator defined as $\hat{s}_{\mu}:=-\hat{\sigma}_{\mu}^{(1)}\otimes\hat{\sigma}_{\mu}^{(2)}\otimes\hat{\sigma}_{\mu}^{(3)}$, i.e. $\hat{H}_{3},\hat{s}_{\mu}]=0$ and that $\mathrm{SU}(2)$ group is spanned by $\{\hat{s}_{\mu}\}_{\mu=x,y,z}$. This symmetry structure yields simultaneous eigenstates as follows:
[…]
@@
Figure 6: The quantum circuit of the time evolution operator $\exp\left(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)$ given by Eq. (21).
@@
$\begin{split}\hat{H}_{\mathrm{eff}}&=\hat{U}_{\mathrm{enc}}\hat{H}_{3}\hat{U}_{\mathrm{enc}}^{\dagger}\\ &=\sqrt{2}(\hat{h}^{(1)}+\hat{h}^{(2)})-(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{x}^{(1)}\otimes\hat{\sigma}_{z}^{(2)}).\end{split}$ (19)
where $\hat{h}^{(i)}=(\hat{\sigma}_{x}^{(i)}+\hat{\sigma}_{z}^{(i)})/\sqrt{2}$ is the Hadamard operator. This yields the following equivalence between the time evolution operators:
@@
Figure 7: The quantum circuit of the single Trotter block for the 99-site system, denoting $\Delta t^{\prime}$ time evolution. The two-qubit unitary gate $\hat{u}$ stays the same as the conventional one shwon in Fig. 2. The red dashed rectangles denote the pair of CNOT gates offsetting with each other. The areas of each Trotter gate $\exp\left(-{\mathrm{i}}\hat{A}_{k}\Delta t^{\prime}\right)$ and $\exp\left(-{\mathrm{i}}\hat{B}_{k}\Delta t^{\prime}\right)$ are separated by the dash-dot lines and highlighted by the characters at the bottom of the figure.
@@
To design an efficient circuit implementation of $\exp\left(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)$, we further expand $\exp\left(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)$ in the following form
@@
$\begin{split}\exp\left(-\mathrm{i}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)&=\exp\left(-{\mathrm{i}}(\hat{h}^{(1)}+\hat{h}^{(2)})\Delta t^{\prime}/\sqrt{2}\right)\\ &\quad\times\exp\left({\mathrm{i}}(\hat{\sigma}_{z}^{(1)}\otimes\hat{\sigma}_{x}^{(2)}+\hat{\sigma}_{x}^{(1)}\otimes\hat{\sigma}_{z}^{(2)})\Delta t^{\prime}\right)\\ &\quad\times\exp\left(-{\mathrm{i}}(\hat{h}^{(1)}+\hat{h}^{(2)})\Delta t^{\prime}/\sqrt{2}\right)\\ &\quad+\mathcal{O}(m^{-3}),\end{split}$ (21)
where the third order error $\mathcal{O}(m^{-3})$ is with a higher than that of the conventional Trotter decomposition, derived from the known decomposition
$\begin{split}\exp\left(\frac{\hat{X}+\hat{Y}}{m}\right)=\exp\left(\frac{\hat{X}}{2m}\right)\exp\left(\frac{\hat{Y}}{m}\right)\exp\left(\frac{\hat{X}}{2m}\right)+\mathcal{O}(m^{-3}).\end{split}$ (22)
Thus, the proposed construction does not crucially affect on the numerical error for the time propagation by applying sufficient number of Trotter iterations $m$. This provides a quantum circuit of the time evolution $\exp\left(-{\mathrm{i}}\hat{H}_{\mathrm{eff}}\Delta t^{\prime}\right)$, represented in Fig. 6. The number of required CNOT gates amounts up to eight for implementing $\exp\left(-{\mathrm{i}}\hat{H}_{\rm eff}\Delta t^{\prime}\right)$ as a quantum circuit.
Figure 8: The process infidelity to the theoretically predicted time evolution without noise. Plots of the conventional Trotter decomposition (Fig. 2) are colored red and those of the proposed decomposition (Fig. 7) are colored blue. (a) The process infidelity of the evolved state to the ideally evolved state without noise, scaling with the number of Trotter iterations. The x-axis $k$ denotes the $k$-th iteration of $n\in\{40,60,80,\ldots,400\}$ for the conventional method and $m\in\{20,30,40,\ldots,200\}$ for the proposed method to make the time intervals $\Delta t$ and $\Delta t^{\prime}$ consistent between each method, i.e. $\Delta t=\Delta t^{\prime}$. (b) The process infidelity of the evolved state to the ideally evolved state without noise, scaling with the number of CNOT gates.